On January 1, 20X1, the Holloran Corporation purchased a machine at a cost of $55,000. The machine was expected to have a service life of 10 years and a $5,000 residual value. The straight-line depreciation method was used. In 20X3, the company switched to the double-declining-balance depreciation method. Depreciation for 20X3 should be:

Multiple Choice

  • $12,750

  • $10,000

  • $11,250

  • $8,500
Answer #1
Annual straight-line depreciation 5000 =(55000-5000)/10
Accumulated Depreciation for 2 years 10000 =5000*2
Book value, January 1, 20X3 45000 =55000-10000
Double-declining-balance depreciation date 25% =1/8*2
Depreciation for 20X3 11250 =45000*25%
Option C $11,250 is correct